OpenStudy (anonymous):

Find the next 4 terms of the sequence. a1= -5, an = an-1-8. a2= a3= a4= a5=

OpenStudy (mathmate):

a1=-5 a2=an-1 -8 = -13 a3=an-2 -8 = -21, ...

OpenStudy (anonymous):

Oh duh! Thanks mathmate! :)

OpenStudy (mathmate):

You're welcome!